### v1

# Reed-Solomon Codes Against Insertions and Deletions: Full-Length and Rate-$1/2$ Codes

Abstract.
The performance of Reed–Solomon codes (RS codes, for short) in the presence of insertion and deletion errors has been studied recently in several papers. In this work, we further study this intriguing mathematical problem, focusing on two regimes. First, we study the question of how well full-length RS codes perform against insertions and deletions. For 2-dimensional RS codes, we fully characterize which codes cannot correct even a single insertion or deletion and show that (almost) all 2-dimensional RS codes correct at least $1$ insertion or deletion error. Moreover, for large enough field size $q$, and for any $k\geq 2$, we show that there exists a full-length $k$-dimensional RS code that corrects $q/10k$ insertions and deletions. Second, we focus on rate $1/2$ RS codes that can correct a single insertion or deletion error. We present a polynomial time algorithm that constructs such codes for $q=O(k^{4})$. This result matches the existential bound given in [1].
R. C. and E. Y. are supported by the European Union (DiDAX, 101115134). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. A. G. and M. M. are supported by the Villum Fonden through grant VIL”52303”.
1. Introduction
This work considers the model of insertions and deletions (insdel errors, for short), the most common model of synchronization errors. An insertion error is when a new symbol is inserted between two adjacent symbols of the transmitted word and a deletion is when a symbol is removed from the transmitted word. Note that this type of errors, unlike substitutions or erasures, can change the length of the transmitted message. Insdel errors cause loss of synchronization between the sender and the receiver, which makes the task of designing codes (over small alphabets) for this model a very tantalizing question.
This natural theoretical model, together with possible application in various fields, including in DNA-Based storage systems, has led many researchers to construct and study codes against insdel errors (Most of the results can be found in the following excellent surveys [2, 3]).
In this work, we focus on one of the most well-known family of linear codes, called Reed–Solomon codes (RS codes, for short), which are defined as follows.
Definition 1 (Reed–Solomon code).
Let $\alpha_{1},\alpha_{2},\ldots,\alpha_{n}$ be distinct elements in the finite field $\mathbb{F}_{q}$ with $q$ elements. For $k<n$, the $[n,k]_{q}$ Reed–Solomon (RS) code of dimension $k$ and block length $n$ associated with the evaluation vector ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{F}_{q}^{n}$ is defined to be the set of codewords
$\text{RS}_{n,k}({\bm{\alpha}}):=\left\{\left(f(\alpha_{1}),\ldots,f(\alpha_{n})\right):f\in\mathbb{F}_{q}[x]_{<k}\right\}.$
RS codes play an important role in ensuring data integrity across various media. Several applications include QR codes, distributed data storage, and data transmission over noisy channels. Furthermore, RS codes have found many theoretical applications in cryptography, and theoretical computer science. The appeal of RS codes is due to their simple algebraic structure, which provides efficient encoding and decoding algorithms, and also shows that they have optimal rate-error trade-off in the Hamming metric. Thus, it is natural to ask how well RS codes perform against insdel errors. This has been studied in several papers [4, 5, 6, 7, 8, 1, 9, 10, 11]. However, many unsolved questions still remain and there is much to discover.
In this paper, we study two regimes of RS codes. First, we focus on full-length RS codes, that is, $q=n$. By characterizing “bad” permutations on the elements of $\mathbb{F}_{q}$, we show that almost all $2$-dimensional RS codes can correct at least $1$ insdel error. Moreover, we show that for $q=2^{O(k)}$, there exists a $[q,k]_{q}$ RS code that can correct $q/(16k)$ insdel errors. Second, we focus on rate $1/2$ RS codes. We give an algorithm that runs in polynomial time in $k$, that constructs a $[2k,k]_{q}$ RS code correcting a single insdel error where $q=O(k^{4})$. We note that the explicit constructions in [1, 10] require a field of size $2^{k^{k}}$ and that our result matches the existential result from [1].
1.1. Previous Work
(a) Non-linear insdel codes
The study of codes that can correct adversarial insertions and deletions (synchronization errors) started with the seminal works of Levenshtein [12]who showed that the codes of Varshamov and Tenengolts [13] (correcting asymmetric errors) are optimal binary codes that can correct from a single insdel error. The quest for constructing (close to) optimal codes that can correct a constant number of insertions or deletions (even just two), spans many works in recent years with some astonishing results [14, 15, 16, 17, 18, 19, 20]. Despite all of this progress, the question of determining the optimal redundancy-error trade-off for codes correcting a constant number of deletions is still open. When it comes to correcting a fraction of insdel errors, Haeupler and Shahrasbi [21] presented an efficient code with rate $1-\delta-\varepsilon$ over an alphabet of size $O_{\varepsilon}(1)$ that can correct from a $\delta$ fraction of insdel errors. These codes are optimal in the sense that they can get as close as we want to the Singleton bound, which is $1-\delta$. For binary codes, the gap between the upper [22] and lower bound [23] on the rate-error trade-off is huge.
(b) Linear insdel codes and RS codes
Reading the above paragraph, one might ask: “How come no code among the referenced ones, is a linear code?”. The reason appears in [24] where it was shown that any linear code correcting a single insdel error must have rate at most $1/2$. This shows that linear codes are provably worse than non-linear as non-linear codes correcting a single insdel error can have rate $1-o(1)$. Then, in [25], the authors generalized this result and proved the half-Singleton bound which states that any $[n,k]_{q}$ linear code can correct at most $n-2k+1$ insdel errors.
In this work, we focus on RS codes which by definition, require the alphabet size $q$ to be at least $n$. As far as we know, the performance of RS codes against insdel errors was first considered in [4] in the context of the traitor tracing problem. In [5, Theorem 3.2], the authors constructed a $[5,2]_{q}$ generalized RS code that can correct a single deletion when $q>8$. They also showed how to extend this construction for any $k$ (by induction), but only provided that there exists a choice of the evaluation points $\alpha_{i}$ and multipliers $v_{i}$ satisfying some specific properties [5, Theorem 3.3]. The rate $1/2$ resulting codes are never usual RS codes. This can be seen in [5, Theorem 3.2] where in fact the choice $v_{i}=1$ for all $i$ is not allowed. In [6], the authors constructed $[n,k]_{q}$, (where $n$ can be as large as $q$) RS codes that can correct from $\log_{k}(q)-1$ insdel errors. However, their construction relies on the existence of a polynomial of degree $k$ that has special properties. The authors do not provide a method for constructing such a polynomial nor do they show its existence (for general $k$ and $q$).
Then, in [1], the authors showed the existence of $[n,k]_{q}$ RS codes that can correct from $n-2k+1$ insdel errors where
$q=O\left(\binom{n}{2k-1}^{2}k^{2}\right).$
These codes attain the half-Singleton bound with equality. They also provided an explicit construction of such codes where $q\approx n^{k^{k}}$. In [11], it was shown that there are $[n,k]_{q}$ RS codes that can correct from $(1-\varepsilon)n-2k+1$ insdel errors where $q=O(n+2^{\textup{poly}(1/\varepsilon)}k)$. In [7, 8, 1, 9], the particular case of $k=2$ was considered and it was shown that the minimal field size of an $[n,2]_{q}$ code that can correct from $n-3$ deletions is $q=\Theta(n^{3})$.
1.2. Our Contribution
In this work, we consider two very basic questions regarding the performance of RS codes against insdel errors.
Question 1.
Can a full-length RS code correct insdel errors, and if it can, how many?
Consider the scenario where our RS code is defined over $\mathbb{F}_{p}$, for a prime $p$, and ${\bm{\alpha}}=(0,1,2,\ldots,p-1)$. Then this $\text{RS}_{n,k}({\bm{\alpha}})$ code over $\mathbb{F}_{p}$ cannot correct even a single deletion. Indeed, if we delete the first symbol from the codeword that corresponds to $f(x)=x$ and the last symbol from the codeword that corresponds to $g(x)=x+1$, then we get the same vector of length $p-1$. However, what happens if we consider “less natural” orderings on the points? Can we choose a different permutation of $\mathbb{F}_{p}$ for which the respective 2-dimensional RS code can correct a single deletion, or even more than that? We answer this question affirmatively. We show that almost all permutations give rise to a 2-dimensional RS code that can correct (at least) a single deletion. Moreover, we show – using a probabilistic argument – that for large enough $q$, there exists an ordering of $\mathbb{F}_{q}$ such that the respective RS code can correct from $q/(10k)$ insdel errors. Formally, we prove the following theorem.
Theorem 1.
Let $k$ be an integer and $q$ be a prime power such that $q\geq e^{9k}\cdot(16ek^{3})$. There exists an evaluation vector ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{q})$ such that the code $\textup{RS}_{q,k}({\bm{\alpha}})$ can correct from any $\frac{q}{16k}$ insdels.
The second question considers rate $1/2$ RS codes that can correct a single deletion. Recall that any linear code correcting a single insdel error must have rate at most $1/2$ and for $k=1$ it is achieved by the trivial repetition code. For larger dimensions it is not obvious what codes of rate $1/2$ can correct a single deletion.
Question 2.
Construct RS codes of rate exactly $1/2$ that can correct $1$ insdel error.
We start by presenting an $[4,2]_{7}$ RS code that can correct a single insdel error and show that $q=7$ is the minimal field size for such a code. Then, this code will be used for an inductive process in which we construct an $[2k,k]_{q}$ RS code correcting a single insdel error. Formally,
Theorem 2.
Let $q=O(k^{4})$ be a prime power. There exists a polynomial time algorithm that outputs ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{2k})\in\mathbb{F}_{q}^{2k}$ for which the respective $\text{RS}_{n,k}({\bm{\alpha}})$ code can correct from a single insdel error.
2. Preliminaries
Throughout, $\mathbb{F}_{q}$ will denote a finite field of order $q$ and $\mathbb{F}_{q}[X_{1},\ldots,X_{n}]$ denotes the ring of polynomials in $X_{1},\dots,X_{n}$ over $\mathbb{F}_{q}$. We recall the notions of a subsequence and a longest common subsequence.
Definition 2.
A subsequence of a string ${\bf s}$ is a string obtained by removing some (possibly none) of the symbols in ${\bf s}$.
Definition 3.
Let ${\bf s},{\bf s}^{\prime}$ be strings over an alphabet $\Sigma$. A longest common subsequence between ${\bf s}$ and ${\bf s}^{\prime}$, is a subsequence of both ${\bf s}$ and ${\bf s}^{\prime}$, of maximal length. We denote by $\textup{LCS}({\bf s},{\bf s}^{\prime})$ the length of a longest common subsequence of ${\bf s}$ and ${\bf s}^{\prime}$.
The edit distance or insdel distance between ${\bf s}$ and ${\bf s}^{\prime}$, denoted by $\textup{ED}({\bf s},{\bf s}^{\prime})$, is the minimal number of insertions and deletions needed in order to turn ${\bf s}$ into ${\bf s}^{\prime}$.
For a code $\mathcal{C}$ we use the following notations:
$\textup{LCS}(\mathcal{C})$ $:=\max\{\textup{LCS}({\bf c},{\bf c}^{\prime}):{\bf c},{\bf c}^{\prime}\in\mathcal{C}\},$
$\textup{ED}(\mathcal{C})$ $:=\min\{\textup{ED}({\bf c},{\bf c}^{\prime}):{\bf c},{\bf c}^{\prime}\in\mathcal{C}\}.$
We have $\textup{ED}(\mathcal{C})=2n-2\textup{LCS}(\mathcal{C})$ if $\mathcal{C}$ has length $n$. It is well known that the insdel correction capability of a code is determined by the LCS of its codewords. Specifically,
Lemma 1.
A code $\mathcal{C}$ can correct $\delta n$ insdel errors if and only $\textup{LCS}({\bf c},{\bf c}^{\prime})\leq n-\delta n-1$ for any distinct ${\bf c},{\bf c}^{\prime}\in\mathcal{C}$, i.e., $\textup{LCS}(\mathcal{C})\leq n-\delta n-1$.
We observe that the edit distance $\textup{ED}(\cdot,\cdot)$ satisfies the following for any ${\bf c},{\bf c}^{\prime}\in\mathbb{F}_{q}^{n}$:
$\textup{ED}({\bf c},{\bf c}^{\prime})$ $=\textup{ED}(\lambda{\bf c},\lambda{\bf c}^{\prime}),\quad\text{for any $\lambda\in\mathbb{F}_{q}^{*}$,}$ (1)
$\textup{ED}({\bf c},{\bf c}^{\prime})$ $=\textup{ED}({\bf c}+\mathbf{1},{\bf c}^{\prime}+\mathbf{1}),$ (2)
where $\mathbf{1}=(1,1,\dots,1)\in\mathbb{F}_{q}^{n}$. We denote the Hamming distance between ${\bf c}$ and ${\bf c}^{\prime}$ by $d_{H}({\bf c},{\bf c}^{\prime})$ and the Hamming-weight of ${\bf c}$ by $w_{H}({\bf c})$.
In this paper we are interested in RS codes and their insertion deletion correction capabilities. We start by citing the (non-asymptotic version of) rate-error–correction trade-off for linear codes correcting insdel errors.
Theorem 3 (Half-Singleton bound; see [26, Corollary 5.2]).
An $[n,k]_{q}$ linear code $\mathcal{C}$ can correct at most $n-2k+1$ insdel errors. Equivalently, $\textup{LCS}(\mathcal{C})\geq 2k-2$.
We call codes attaining the bound of Theorem 3 optimal codes in this paper.
Notation 1.
We say that a vector of indices $I\in[n]^{\ell}$ is an increasing sequence if its coordinates are monotonically strictly increasing, i.e., for any $1\leq i<j\leq\ell$, $I_{i}<I_{j}$, where $I_{i}$ is the $i$th coordinate of $I$. For an increasing vector $I\in[n]^{\ell}$ and an evaluation vector ${\bm{\alpha}}=(\alpha_{1},\dots,\alpha_{n})\in\mathbb{F}_{q}^{n}$, we denote by ${\bm{\alpha}}_{I}$ the subsequence of ${\bm{\alpha}}$ indexed by $I$, i.e., ${\bm{\alpha}}_{I}:=(\alpha_{I_{1}},\dots,\alpha_{I_{\ell}})$. Moreover, for $f\in\mathbb{F}_{q}[x]$ we let $f({\bm{\alpha}}_{I}):=(f(\alpha_{I_{1}}),\dots,f(\alpha_{I_{\ell}}))$.
For two increasing sequences $I,J\in[n]^{\ell}$, we define the following matrix of order ×$\ell\times(2k-1)$ in the formal variables ${\bf X}=(X_{1},\ldots,X_{n})$, which we denote by $V_{k,\ell,I,J}({\bf X})$:
$V_{k,\ell,I,J}({\bf X})=\begin{pmatrix}1&X_{I_{1}}&\ldots&X_{I_{1}}^{k-1}&X_{J_{1}}&\ldots&X_{J_{1}}^{k-1}\\ 1&X_{I_{2}}&\ldots&X_{I_{2}}^{k-1}&X_{J_{2}}&\ldots&X_{J_{2}}^{k-1}\\ \vdots&\vdots&\ldots&\vdots&\vdots&\ldots&\vdots\\ 1&X_{I_{\ell}}&\ldots&X_{I_{\ell}}^{k-1}&X_{J_{\ell}}&\ldots&X_{J_{\ell}}^{k-1}\\ \end{pmatrix}.$ (3)
In [11], the following algebraic condition was proved.
Lemma 2 (see [11, Lemma 12]).
If for increasing sequences $I,J\in[n]^{2k-1}$ with $d_{H}(I,J)\geq k$, it holds that $\det(V_{k,2k-1,I,J}({\bm{\alpha}}))\neq 0$, then the code $\text{RS}_{n,k}({\bm{\alpha}})$ is optimal. Moreover, if the code is optimal, then the only vectors in $\text{Ker}\left(V_{k,2k-1,I,J}({\bm{\alpha}})\right)$ are of the form $(0,f_{1},\ldots,f_{k-1},-f_{1},\ldots,-f_{k-1})$ for some $f_{i}\in\mathbb{F}_{q}$.
3. Full-Length RS Codes
In this section we consider full-length RS codes ($q=n$), i.e., the evaluation vector ${\bm{\alpha}}=(\alpha_{1},\dots,\alpha_{q})$ is such that $\{\alpha_{1},\dots,\alpha_{q}\}=\mathbb{F}_{q}$. We show how the order in which the elements of $\mathbb{F}_{q}$ show up in ${\bm{\alpha}}$ heavily influences the insdel distance of the code.
3.1. $k=2$: Correcting a Single Insdel Error
Definition 4.
We say that ${\bm{\alpha}}=(\alpha_{1},\dots,\alpha_{q})\in\mathbb{F}_{q}^{q}$ and $\widetilde{{\bm{\alpha}}}=(\widetilde{\alpha}_{1},\dots,\widetilde{\alpha}_{q})\in\mathbb{F}_{q}^{q}$ are equivalent if $(\lambda\alpha_{1}+\mu,\dots,\lambda\alpha_{q}+\mu)=\widetilde{{\bm{\alpha}}}$ for some $\mu\in\mathbb{F}_{q}$ and $\lambda\in\mathbb{F}_{q}^{*}$.
It is easy to see that for equivalent ${\bm{\alpha}},\widetilde{{\bm{\alpha}}}\in\mathbb{F}_{q}^{n}$ we have $\text{RS}_{n,2}({\bm{\alpha}})=\text{RS}_{n,2}(\widetilde{{\bm{\alpha}}})$. Moreover, in every equivalence class, except for the case where ${\bm{\alpha}}$ is a constant vector, there are $q(q-1)$ vectors (we have $q$ choices for $\mu$ and $q-1$ choices for $\lambda$).
Lemma 3.
For be a primitive element $\theta$ of $\mathbb{F}_{q}$, consider the following vectors:
  1. (1)
    $(0,1,\theta,\dots,\theta^{q-2})$;
  2. (2)
    $(\theta^{q-2},\dots,\theta,1,0)$;
  3. (3)
    $(0,1,2,\dots,q-1)$ if $q$ is prime.
Then, $\text{RS}_{q,2}({\bm{\alpha}})$ cannot correct a single insdel error if and only if ${\bm{\alpha}}$ is equivalent to one of the vectors in 1)3)1)-3)1 ) - 3 ).
Proof.
Suppose that ${\bm{\alpha}}$ is such that $\text{RS}_{q,2}({\bm{\alpha}})$ has insdel distance 2. Then there exist increasing sequences $I,J\in[q]^{q-1}$ with $d_{H}(I,J)\geq(q-1)-1=q-2$, and with the property that $f_{a,b}({\bm{\alpha}}_{I})={{\bm{\alpha}}}_{J}$ for some $f_{a,b}\in AGL(\mathbb{F}_{q})$; see Notation 2. We start with the case $d_{H}(I,J)=q-1$. The only possibilities for $I,J\in[q]^{q-1}$ are then $(2,\dots,q)$ and $(1,\dots,q-1)$. W.l.o.g. let $I=(2,\dots,q)$ and $J=(1,\dots,q-1)$. By subtracting an appropriate constant vector and multiplying by a scalar, and because equivalent vectors give rise to the same code, we can further assume that ${\bm{\alpha}}=(0,1,\alpha_{3},\dots,\alpha_{q})$. Since $f_{a,b}({\bm{\alpha}}_{I})={{\bm{\alpha}}}_{J}$ we obtain the following system of equations:
$\begin{cases}a\cdot 0+b&=1\\ a\cdot 1+b&=\alpha_{3}\\ a\cdot\alpha_{3}+b&=\alpha_{4}\\ &\vdots\\ a\cdot\alpha_{q-1}+b&=\alpha_{q}\\ a\cdot\alpha_{q}+b&=0\end{cases}$
where the last equation comes from the fact that $f_{a,b}$ acts bijectively on $\mathbb{F}_{q}$. From the above set of equations we obtain
$b=1,\quad\alpha_{3}=a+1,\quad\dots,\quad\alpha_{q}=a^{q-2}+\dots+a+1.$
Since additionally we know that $a\cdot\alpha_{q}+1=0$, we have that $a^{q-1}+\dots+a+1=0$. If $a\neq 1$, then $(a^{q-1}+\dots+a+1)=\frac{a^{q}-1}{a-1}=0$. Therefore $a^{q}=1$ implying $a=1$, contradiction. Thus the only solution is $a=1$, and from the above set of equations we obtain that ${\bm{\alpha}}=(0,1,2,\dots,q-1)$.
Now suppose that $d_{H}(I,J)=q-2$. This implies that either $I$ and $J$ are equal to $(2,\dots,q)$ and $(1,\dots,q-2,q)$ or to $(1,3,\dots,q)$ and $(1,\dots,q-1)$. Assume that $I=(2,\dots,q)$ and $J=(1,\dots,q-2,q)$. As before, through scaling, we can assume that ${\bm{\alpha}}=(\alpha_{1},\dots,\alpha_{q-2},1,0)$. Since $f_{a,b}({\bm{\alpha}}_{I})={{\bm{\alpha}}}_{J}$ we obtain:
$\begin{cases}a\cdot 0+b&=0\\ a\cdot 1+b&=\alpha_{q-3}\\ a\cdot\alpha_{q-3}+b&=\alpha_{q-4}\\ &\vdots\\ a\cdot\alpha_{2}+b&=\alpha_{1}\\ a\cdot\alpha_{1}+b&=1\end{cases}$
where again the last equation comes from the fact that $f_{a,b}$ is a bijection. We have
$b=0,\quad\alpha_{q-3}=a,\quad,\alpha_{q-4}=a^{2},\quad\dots,\quad\alpha_{1}=a^{q-2}.$
Since $a\cdot\alpha_{1}+b=1$ we also have $a^{q-1}=1$. Since we assume that ${\bm{\alpha}}$ is an evaluation vector for an RS code over $\mathbb{F}_{q}$ of length $q$, we need to have $\{\alpha_{1},\dots,\alpha_{q}\}=\{0,1,a,\dots,a^{q-2}\}=\mathbb{F}_{q}$ and thus $a$ is a primitive element of $\mathbb{F}_{q}$. Therefore ${\bm{\alpha}}$ is equivalent to $(a^{q-2},\dots,a^{2},a,1,0)$.
The case where $I=(1,3,\dots,q)$ and $J=(1,\dots,q-1)$ can be proven analogously, where one arrives at the conclusion that ${\bm{\alpha}}$ is equivalent to $(0,1,a,a^{2},\dots,a^{q-2})$ for some primitive element $a$ of $\mathbb{F}_{q}$.
Now suppose that ${\bm{\alpha}}$ is equivalent to one of the vectors in $(1)-(3)$. It is enough to show that for any of the vectors listed in this lemma, the 2-dimensional RS code with that vector as the evaluation vector ${\bm{\alpha}}$, contains a codeword, ${\bf c}$ different from the evaluation vector, which shares a common subsequence of length $q-1$ with the evaluation vector.
  1. (1)
    If ${\bm{\alpha}}=(0,1,\theta,\dots,\theta^{q-2})$, then let ${\bf c}=\theta{\bm{\alpha}}=(0,\theta,\dots,\theta^{q-2},1)$. Therefore we have $\text{LCS}({\bm{\alpha}},{\bf c})=q-1$.
  2. (2)
    If ${\bm{\alpha}}=(\theta^{q-2},\dots,\theta,1,0)$, then let ${\bf c}=\theta{\bm{\alpha}}=(1,\theta^{q-2},\dots,\theta,0)$. Therefore we have $\text{LCS}({\bm{\alpha}},{\bf c})=q-1$.
  3. (3)
    If ${\bm{\alpha}}=(0,1,\dots,q-1)$, then let ${\bf c}={\bm{\alpha}}+\textbf{1}=(1,2,\dots,q-1,0)$. Therefore we have $\text{LCS}({\bm{\alpha}},{\bf c})=q-1$. ∎
In the next proposition, which is a consequence of Lemma 3, $\phi$ denotes Euler’s totient function, i.e., $\phi(n)$ is the number of positive integers up to $n$ that are relatively prime to $n$.
Proposition 1.
There are at least
$\begin{cases}(q-2)!-2\phi(q-1)-1&\textnormal{ if $q$ is prime,}\\ (q-2)!-2\phi(q-1)&\textnormal{ if $q$ is not prime,}\end{cases}$
orderings ${\bm{\alpha}}=(\alpha_{1},\dots,\alpha_{q})$ of the elements in $\mathbb{F}_{q}$ such that $\text{RS}_{q,2}({\bm{\alpha}})$ can correct at least a single insdel error.
Proof.
From Lemma 3 we know the vectors that give rise to codes with insdel distance 2. Therefore, for every primitive element of $\mathbb{F}_{q}$ we have 2 equivalence classes, and if $q$ is prime we also have the equivalence class represented by $(0,1,2,\dots,q-1)$. In total there are $(q-2)!$ equivalence classes, which gives the statement of the lemma. ∎
Some explicit evaluations for the bound of Proposition 1 are the following.
$q$ 4 5 7 8 9 11 13
prop. 0.000 0.333 0.967 0.983 0.998 0.999 0.999
Table 1. Proportion of 2-dimensional full-length RS codes correcting at least a single insdel error within the set of all codes for various values of prime powers $q$.
It is well-known that $\sqrt{n/2}\leq\phi(n)\leq n-1$. From this it follows that when $q$ is large, then almost all (meaning the proportion of such codes approaches 1 as $q\to+\infty$) full-length 2-dimensional RS codes can correct a single insdel error. More precisely, we have the following result.
Corollary 1.
We have
$\lim_{q\to+\infty}\frac{|\{{\bm{\alpha}}\in\mathbb{F}_{q}^{q}:\textnormal{$\textup{RS}_{q,2}({\bm{\alpha}})$ can correct at least a single insdel error}\}|}{|\{{\bm{\alpha}}\in\mathbb{F}_{q}^{q}\}|}=1.$
3.2. General $k$
Our next goal is to use the power of randomness and show that there is a positive probability that a random permutation of $\mathbb{F}_{q}$ (when $q$ is large enough compared to $k$) will give rise to an RS code that can correct many insdel errors. In the following, we prove Theorem 1 which is restated for convenience. See 1 We will need the following claim in the proof.
Claim 1.
Let ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{q})\in\mathbb{F}_{q}^{q}$ be a uniformly random vector. Then, the probability that all distinct $i,j\in[q]$, we have $\alpha_{i}\neq\alpha_{j}$ is at least $e^{-q}$.
Proof.
The number of permutations on $q$ elements is $q!$. Thus, this probability is exactly $q!/q^{q}$. By Stirling’s approximation (e.g., [27, Lemma 7.3]), we get
$\frac{q!}{q^{q}}\geq\frac{\sqrt{8\pi q}\left(\frac{q}{e}\right)^{q}}{q^{q}}\geq e^{-q}\;.\qed$
Proof of Theorem 1.
Let ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{q})\in\mathbb{F}_{q}^{q}$ be a uniform random vector. Note that it can be that ${\bm{\alpha}}$ is not a permutation of $\mathbb{F}_{q}$ and we will address this issue at the end of the proof. Set $\ell=q-q/(16k)$. For two increasing sequences $I,J\subset[q]^{\ell}$ that agree on at most $k-1$ coordinates, we will bound from above the probability that $V_{k,\ell,I,J}({\bm{\alpha}})$ is not full rank.
For this purpose, consider the matrix $V_{k,\ell,I,J}({\bf X})$ from (3) where ${\bf X}=(X_{1},\ldots,X_{n})$ are formal variables. Consider the matrix $M({\bf X})$ obtained by taking the top $(2k-1)\times(2k-1)$ submatrix of $V_{k,\ell,I,J}({\bf X})$. Its determinant is a nonzero polynomial [1, Proposition 18] of degree less than $k^{2}$. Thus, by the Schwartz-Zippel Lemma [28, 29], $\Pr[\det(M({\bm{\alpha}}))=0]\leq k^{2}/q$ where the probability is over all ${\bm{\alpha}}\in\mathbb{F}_{q}^{q}$.
We shall construct a sequence $M_{1}({\bf X}),\ldots,M_{m}({\bf X})$ of $(2k-1)\times(2k-1)$ submatrices of $V_{k,\ell,I,J}({\bf X})$. Denote by $\textup{Vars}(M_{j}({\bf X}))=\{i\in[n]:X_{i}\textup{ appears in }M_{j}({\bf X})\}$. We require our sequence of submatrices to comply with the following condition: for all distinct $i,j\in[m]$ it holds that
$\textup{Vars}(M_{i}({\bf X}))\cap\textup{Vars}(M_{j}({\bf X}))=\emptyset\;.$
In words, this condition assures that for any distinct $i,j\in[m]$, the set of variables that appear in $M_{i}$ is disjoint from the set of variables that appear in $M_{j}$. This implies that the event $\det(M_{i}({\bm{\alpha}}))=0$ is independent of all the events $\{\det(M_{j}({\bm{\alpha}}))=0:j\in[m]\setminus\{i\}\}$ for all $i\in[m]$. Therefore,
$\Pr_{{\bm{\alpha}}\sim\mathbb{F}_{q}^{q}}[\textup{rank}(V_{k,\ell,I,J}({\bm{\alpha}}))<2k-1]$
$\;\;\;\leq\Pr_{{\bm{\alpha}}\sim\mathbb{F}_{q}^{q}}[\forall i\in[m],\;\det(M_{i}({\bm{\alpha}}))=0]$
$\;\;\;\leq\prod_{i\in[m]}\Pr_{{\bm{\alpha}}\sim\mathbb{F}_{q}^{q}}[\det(M_{i}({\bm{\alpha}})=0)]\leq\left(\frac{k^{2}}{q}\right)^{m}\;,$
where the first inequality is by observing that if $\textup{rank}(V_{k,\ell,I,J}({\bm{\alpha}}))({\bm{\alpha}})<2k-1$ then each $(2k-1)\times(2k-1)$ submatrix must be singular and because of the independence argument above, the probabilities for each matrix can be multiplied.
Next, we claim that $m$ is at least $\ell/6k$. Indeed, if one wishes to greedily construct such a sequence of matrices then the first matrix $M_{1}$ can be chosen by picking the first $12k-$ 1 rows out of the $\ell$ rows. Next, we need to choose $M_{2}$ such that $\textup{Vars}(M_{1}({\bf X}))\cap\textup{Vars}(M_{2}({\bf X}))=\emptyset$. By the definition of $V_{k,\ell,I,J}({\bf X})$ (see (3)), every variable can appear at most twice. Then, the number of rows from $V_{k,\ell,I,J}({\bf X})$ that can be taken to form $M_{2}$ is at least $\ell-3(2k-1)$ (the rows of $M_{1}$ also cannot be chosen again). Thus, continuing this way for $M_{3},\ldots,M_{m}$, we get that $m\geq\ell/(6k)$.
Thus, for fixed increasing sequences $I,J$ of length $\ell$ that agree on at most $k-1$ coordinates, the probability that $V_{k,\ell,I,J}({\bm{\alpha}})$ is not full rank is at most
$\left(\frac{k^{2}}{q}\right)^{\ell/(6k)}\;.$
By the union bound, the probability that there exists a pair $I,J$ of increasing sequences of length $\ell$ that agree on at most $k-1$ coordinates for which the rank of $V_{k,\ell,I,J}({\bm{\alpha}})$ is not full, is at most
$\left(\frac{k^{2}}{q}\right)^{\ell/(6k)}\binom{q}{\ell}^{2}$ $\leq\left(\frac{k^{2}}{q}\right)^{q/(8k)}\left(\frac{eq}{\frac{q}{16k}}\right)^{q/(8k)}$
$=\left(\frac{16ek^{3}}{q}\right)^{q/(8k)}\leq e^{-\frac{9}{8}q},$
where the first inequality is by our requirement that the code can correct $q/(16k)$ insdel errors, which implies that $\ell=q-q/(16k)$ and then for any integer $q\geq 64k$, we have that $\ell/(6k)>\ell/(8k)$. Furthermore, we used the inequality $\binom{n}{k}<(en/k)^{k}$. The final inequality follows by our assumption that $q\geq e^{9k}\cdot(16ek^{3})$.
Finally, observe that in order for our randomly chosen ${\bm{\alpha}}$ to define a proper RS code, it must be that all points in ${\bm{\alpha}}$ are distinct. From Claim 1 we know that the probability that this happens is at least $e^{-q}$. Thus, a random vector ${\bm{\alpha}}\in\mathbb{F}_{q}^{q}$ does not give rise to an RS code that can correct from $q/(16k)$ insdel errors if ${\bm{\alpha}}$ is not a permutation or it is a permutation but there exists a pair of increasing sequences $I,J\subset[n]^{q-q/(16k)}$ that agree on at most $k-1$ coordinates such that $\det(V_{k,q-q/(16k),I,J}({\bm{\alpha}}))=0$. This occurs with probability at most $(1-e^{-q})+e^{-9q/8}<1$. We conclude that there exists an ${\bm{\alpha}}\in\mathbb{F}_{q}^{n}$ such that the respective $\text{RS}_{q,k}(\alpha_{1},\ldots,\alpha_{q})$ can correct $q/(16k)$ insdel errors. ∎
4. Existence Results
In this section, we give new existence results for optimal RS codes. Our approach is based on the following (easy) lemma, which can be seen as a generalization of [30, Lemma 4.10] for general $k$.
Lemma 4.
A code $\text{RS}_{n,k}({\bm{\alpha}})$ is optimal if and only if $f({\bm{\alpha}}_{I})\neq{g}({\bm{\alpha}}_{J})$ for any $f\in\{f_{k-1}x^{k-1}+\dots+f_{1}x:f_{i}\in\mathbb{F}_{q}\textnormal{ for all $i\in\{1,\dots,k-2\}$},f_{k-1}\in\{0,1\}\}$ and $g\in\mathbb{F}_{q}[x]_{<k}$, with $f\neq g$ and increasing sequences $I,J\in[n]^{2k-1}$.
Proof.
Suppose there exist $f(x)\in\{f_{k-1}x^{k-1}+\dots+f_{1}x:f_{i}\in\mathbb{F}_{q}\textnormal{ for all $i\in\{1,\dots,k-2\}$},$ $f_{k-1}\in\{0,1\}\}$ and $g(x)\in\mathbb{F}_{q}[x]_{<k}$ with $f\neq g$ and increasing vectors $I,J\in[n]^{2k-1}$ such that $I^{f}_{\bm{\alpha}}=J^{g}_{\bm{\alpha}}$. Since $f\neq g$, $f({\bm{\alpha}})\neq g({\bm{\alpha}})$ and $\textup{LCS}(f({\bm{\alpha}}),g({\bm{\alpha}}))\geq 2k-1$ because they at least share the subsequence indexed by $I$ and $J$, respectively. Thus the code is not optimal.
Now suppose that the code $\text{RS}_{n,k}({\bm{\alpha}})$ satisfies $\textup{LCS}(\text{RS}_{n,k}({\bm{\alpha}}))\geq 2k-1$. Then there exist two distinct codewords ${\bf c},{\bf c}^{\prime}\in\textup{RS}_{n,k}({\bm{\alpha}})$ with $\textup{LCS}({\bf c},{\bf c}^{\prime})\geq 2k-1$. W.l.o.g. we assume $\textup{LCS}({\bf c},{\bf c}^{\prime})=2k-1$. We can write ${\bf c}=f({\bm{\alpha}})$ and ${\bf c}^{\prime}=g({\bm{\alpha}})$ for some distinct $f,g\in\mathbb{F}_{q}[x]_{<k}$. Denote the indices of the largest common subsequence of ${\bf c}$ and ${\bf c}^{\prime}$ (i.e. the coordinates where they coincide) by $I$ and $J$, respectively. We have $I,J\in[n]^{2k-1}$ and $f({\bm{\alpha}}_{I})=g({\bm{\alpha}}_{J})$. By the equations in (1) and (2), we can subtract and multiply by suitable elements of $\mathbb{F}_{q}$ to obtain $\textup{LCS}({\bf c},{\bf c}^{\prime})=\textup{LCS}(f({\bm{\alpha}}),g({\bm{\alpha}}))=\textup{LCS}(\tilde{f}({\bm{\alpha}}),\tilde{g}({\bm{\alpha}}))\geq 2k-1$ where $\tilde{f}(x)\in\{\tilde{f}_{k-1}x^{k-1}+\dots+\tilde{f}_{1}x:\tilde{f}_{i}\in\mathbb{F}_{q}\textnormal{ for all $i\in\{1,\dots,k-2\}$},\tilde{f}_{k-1}\in\{0,1\}\}$ and $\tilde{g}\in\mathbb{F}_{q}[x]_{<k}$. ∎
4.1. The Dimension $2$ Case
Already for $k=2$ it is an open problem to determine the smallest $q$ for which there exists an optimal RS code $\text{RS}_{n,2}({\bm{\alpha}})$. In this subsection we prove that $\mathbb{F}_{7}$ is the smallest finite field for which optimal RS codes with these parameters exist. In Subsection 4.2, we build on the result of this subsection and establish the existence of optimal RS codes with rate $1/2$ through induction on $k\geq 2$.
We will need the following notation and lemma, which can be seen as a special case of Lemma 4.
Notation 2.
$AGL(\mathbb{F}_{q})$ denotes the group of affine maps $f_{a,b}$ with $f_{a,b}(x)=ax+b$, $a\in\mathbb{F}_{q}^{*}$ and $b\in\mathbb{F}_{q}$. In other words, given $x\in\mathbb{F}_{q}$, $f_{a,b}(x)=ax+b\in\mathbb{F}_{q}$.
Lemma 5.
(see also [30, Lemma 4.10]) $\text{RS}_{n,2}({\bm{\alpha}})$ has insdel distance $42n-$ 4 if and only if $f_{a,b}({\bm{\alpha}}_{I})={{\bm{\alpha}}}_{J}$ for any $f_{a,b}\in AGL(\mathbb{F}_{q})$ and any two increasing vectors $I,J\in[n]^{3}$ with $d_{H}(I,J)\geq 2$.
Lemma 6.
The smallest $q$ for which there exists an optimal $\text{RS}_{4,2}({\bm{\alpha}})$ code over $\mathbb{F}_{q}$ is $q=7$. In this case, a possible evaluation vector is ${\bm{\alpha}}=(0,1,2,5)\in\mathbb{F}_{7}^{4}$.
Proof.
Because of Lemma 5, the capability of ${\bm{\alpha}}$ giving rise to an optimal $\text{RS}_{4,2}({\bm{\alpha}})$-code over $\mathbb{F}_{q}$ depends on the action of $AGL(\mathbb{F}_{q})$, which is $2$-transitive on $\mathbb{F}_{q}$, see [30, Lemma 2.2 (i)]. This allows to choose the first two coordinates in ${\bm{\alpha}}$ arbitrarily, as long as they are distinct, and so w.l.o.g. we assume that ${\bm{\alpha}}=(0,1,\alpha_{1},\alpha_{2})$ with $\alpha_{1},\alpha_{2}\neq 0,1$ and $\alpha_{1}\neq\alpha_{2}$. From Lemma 5 the code $\text{RS}_{4,2}({\bm{\alpha}})$ has insdel distance $42n-$ 4 if and only if there is no non-trivial $f_{a,b}\in AGL(\mathbb{F}_{q})$ that maps any triple in the following set to any other triple in the same set: $S:=\{(0,1,\alpha_{1}),(0,1,\alpha_{2}),(0,\alpha_{1},\alpha_{2}),(1,\alpha_{1},\alpha_{2})\}$. The strategy now is to compute all the images through $f_{a,b}$ of all the triples in $S$, and force the condition described by the if and only if to be satisfied. Clearly we will assume that $(a,b)\neq(1,0)$, as otherwise $f_{a,b}$ would be the identity map.
  • (i)
    $f_{a,b}(0,1,\alpha_{1})=(b,a+b,a\alpha_{1}+b)\in S$ forces either $b=0$ and $a=\alpha_{1}$ (note that if $a=1$ then $f_{a,b}$ is just the identity map) or $b=1$ and $a=\alpha_{1}-1$. In the first case we get that $f_{a,b}(0,1,\alpha_{1})=(0,\alpha_{1},\alpha_{1}^{2})\in S$ if and only if $\alpha_{2}=\alpha_{1}^{2}$. In the latter case $f_{a,b}(0,1,\alpha_{1})=(1,\alpha_{1},\alpha_{1}^{2}-\alpha_{1}+1)\in S$ if and only if $\alpha_{2}=\alpha_{1}^{2}-\alpha_{1}+1$. In conclusion for all $(a,b)\neq(0,1)$ we have that $f_{a,b}(0,1,\alpha_{1})\not\in S$ exactly when $\alpha_{2}\neq\alpha_{1}^{2},\alpha_{1}^{2}-\alpha_{1}+1$.
  • (ii)
    $f_{a,b}(0,1,\alpha_{2})=(b,a+b,a\alpha_{2}+b)\in S$. As before this condition forces either $b=0$ and $a=\alpha_{1}$ or $b=1$ and $a=\alpha_{1}-1$. In the first case we get that $f_{a,b}(0,1,\alpha_{2})=(0,\alpha_{1},\alpha_{1}\alpha_{2})\in S$ if and only if $\alpha_{1}=1$, which clearly cannot happen. In the latter case $f_{a,b}(0,1,\alpha_{1})=(1,\alpha_{1},\alpha_{2}(\alpha_{1}-1)+1)\in S$ if and only if $\alpha_{2}=\alpha_{2}(\alpha_{1}-1)+1$. In conclusion, for all $(a,b)\neq(0,1)$ we have that $f_{a,b}(0,1,\alpha_{2})\not\in S$ when $\alpha_{2}(\alpha_{1}-2)\neq-1$.
  • (iii)
    $f_{a,b}(0,\alpha_{1},\alpha_{2})=(b,a\alpha_{1}+b,a\alpha_{2}+b)\in S$ forces either $b=0$ and $a=\alpha_{1}^{-1}$ or $b=1$ and $a\alpha_{1}+1=\alpha_{1}$. In the first case we get $f_{a,b}(0,\alpha_{1},\alpha_{2})=(0,1,\alpha_{2}/\alpha_{1})\in S$ if and only if $\alpha_{2}/\alpha_{1}=\alpha_{1}$ or $\alpha_{2}/\alpha_{1}=\alpha_{2}$. This can happen only if $\alpha_{2}=\alpha_{1}^{2}$ as $\alpha_{1}\neq 1$. In the second case we get $f_{a,b}(1,\alpha_{1},\alpha_{2})=(1,\alpha_{1},\alpha_{2}(\alpha_{1}-1)/\alpha_{1}+1)\in S$ if and only if $\alpha_{2}(\alpha_{1}-1)/\alpha_{1}+1=\alpha_{2}$. This case never happens as $\alpha_{1}\neq\alpha_{2}$. In conclusion, for all $(a,b)\neq(0,1)$ we have that $f_{a,b}(0,\alpha_{1},\alpha_{2})=(b,a\alpha_{1}+b,a\alpha_{2}+b)\not\in S$ when $\alpha_{2}\neq\alpha_{1}^{2}$.
  • (iv)
    $f_{a,b}(1,\alpha_{1},\alpha_{2})=(a+b,a\alpha_{1}+b,a\alpha_{2}+b)\in S$ forces $a+b=0$ and either $a\alpha_{1}-a=1$ or $a\alpha_{1}-a=\alpha_{1}$. This is because this time if $a+b=1$ then we need $a\alpha_{1}+b=a\alpha_{1}+1-a=\alpha_{1}$ and hence $a=1$ (recall that $\alpha_{1}\neq 1$). As before, if $(a,b)=(1,0)$ then we just have the identity map and we discard this case. If $a+b=0$ and $a\alpha_{1}-a=1$ then we get $f_{a,b}(1,\alpha_{1},\alpha_{2})=(0,1,(\alpha_{2}-1)/(\alpha_{1}-1))\in S$ only if either $\alpha_{2}=\alpha_{1}^{2}-\alpha_{1}+1$ or $\alpha_{2}(\alpha_{1}-2)=-1$. If $a+b=0$ and $a\alpha_{1}-a=\alpha_{1}$ then $f_{a,b}(1,\alpha_{1},\alpha_{2})=(0,\alpha_{1},(\alpha_{2}-1)\alpha_{1}/(\alpha_{1}-1))$ which is never in $S$ as $\alpha_{1}\neq\alpha_{2}$. In conclusion for all $(a,b)\neq(1,0)$, one has $f_{a,b}(1,\alpha_{1},\alpha_{2})=(a+b,a\alpha_{1}+b,a\alpha_{2}+b)\not\in S$ when $\alpha_{2}\neq\alpha_{1}^{2}-\alpha_{1}+1$ and $\alpha_{2}(\alpha_{1}-2)\neq-1$.
Summarizing up all the cases above we get that $\text{RS}_{4,2}({\bm{\alpha}})$ with ${\bm{\alpha}}=(0,1,\alpha_{1},\alpha_{2})$, $\alpha_{1}\neq 0,1$ and $\alpha_{2}\neq 0,1,\alpha_{1}$ has insertion deletion distance $42n-$ 4 if and only if $\alpha_{2}\not\in\{0,1,\alpha_{1},\alpha_{1}^{2},\alpha_{1}^{2}-\alpha_{1}+1\}$ and $\alpha_{2}\neq-1/(\alpha_{1}-2)$ for $\alpha_{1}\neq 2$. This means $q\geq 7$ is necessary to find a good ${\bm{\alpha}}$, as if $q=4,5$ then there is no $\alpha_{2}\in\mathbb{F}_{q}$ that can satisfy the conditions above for any choice of $\alpha_{1}$. On the other hand if $q=7$ then ${\bm{\alpha}}=(0,1,2,5)$ satisfies all the required conditions. ∎
The following is a consequence of the proof of the previous lemma.
Remark 1.
$\text{RS}_{4,2}({\bm{\alpha}})$ with ${\bm{\alpha}}=(0,1,\alpha_{1},\alpha_{2})$, $\alpha_{1}\neq 0,1$ and $\alpha_{2}\neq 0,1,\alpha_{1}$ is optimal if and only if $\alpha_{2}\not\in\{0,1,\alpha_{1},\alpha_{1}^{2},\alpha_{1}^{2}-\alpha_{1}+1\}$ and $\alpha_{2}\neq-1/(\alpha_{1}-2)$ for $\alpha_{1}\neq 2$.
4.2. Induction on $k$: Rate $1/2$
In this subsection we use the result of the previous subsection as a base case, and we apply induction on $k$ for rate $1/2$ RS codes. We need two claims for proving the main statement which is Proposition 2.
Claim 2.
Let $\ell\geq 2$ and $I,J\in[\ell]^{\ell-1}$ be increasing sequences. Then if $I\neq J$ we have $d_{H}(I,J)=|s_{J}-s_{I}|$ where $s_{I}$ is the unique element in $\{1,\dots,\ell\}\setminus\{I_{1},\dots,I_{(\ell-1)}\}$ and $s_{J}$ is the unique element in $\{1,\dots,\ell\}\setminus\{J_{1},\dots,J_{(\ell-1)}\}$.
Proof.
Suppose $s_{I}=s_{J}$. Then we clearly have $I=J$, contradiction. Therefore we can assume w.l.o.g. that $s_{I}<s_{J}$. We have
$I_{u}=\begin{cases}u&\textnormal{ if $1\leq u<s_{I}$}\\ u+1&\textnormal{ if $s_{I}\leq u\leq\ell-1$}\\ \end{cases}$ (4)
$J_{u}=\begin{cases}u&\textnormal{ if $1\leq u<s_{J}$}\\ u+1&\textnormal{ if $s_{J}\leq u\leq\ell-1$}.\\ \end{cases}$
Therefore we obtain that
$d_{H}(I,J)$ =|$=|\{u\in\{1,\dots,\ell-1\}:I_{u}\neq J_{u}\}|$
$=\ell-1-|\{u\in\{1,\dots,\ell-1\}:I_{u}=J_{u}\}|$
$=\ell-1-(\min\{s_{I},s_{J}\}-1+\ell-\max\{s_{I},s_{J}\})$
$=\ell-1-s_{I}+1-\ell+s_{J}$
$=s_{J}-s_{I}$
which is the statement of the lemma. ∎
Claim 3.
Let $\ell\geq 2$, let ${\bm{\alpha}}=(\alpha_{1},\dots,\alpha_{\ell})\in\mathbb{F}_{q}^{\ell}$ be a vector of pairwise distinct elements of $\mathbb{F}_{q}$ and let $f\in\mathbb{F}_{q}[x]_{<k}$. If for increasing sequences $I,J\in[\ell]^{\ell-1}$ with $d_{H}(I,J)\geq k-1$ we have $f({\bm{\alpha}}_{I})=f({\bm{\alpha}}_{J})$, then $f$ is constant.
Proof.
Let $s_{I}$ and $s_{J}$ be defined as in Claim 2. W.l.o.g. assume that $s_{I}<s_{J}$. Since $f({\bm{\alpha}}_{I})=f({\bm{\alpha}}_{J})$ we have that
$(f(\alpha_{I_{1}}),\dots,f(\alpha_{I_{\ell-1}}))=(f(\alpha_{J_{1}}),\dots,f(\alpha_{J_{\ell-1}}))$
and by the Equations (4) from the proof of Claim 2 we have
$\begin{cases}f(\alpha_{s_{I}+1})&=f(\alpha_{s_{I}}),\\ f(\alpha_{s_{I}+2})&=f(\alpha_{s_{I}+1}),\\ &\vdots\\ f(\alpha_{s_{I}+(s_{J}-s_{I})})&=f(\alpha_{s_{I}+(s_{J}-s_{I})-1}).\end{cases}$
This means in particular that
$f(\alpha_{s_{I}})=f(\alpha_{s_{I}+1})=\dots=f(\alpha_{s_{I}+(s_{J}-s_{I})}),$
where $s_{J}-s_{I}+1=d_{H}(I,J)+1\geq k$. Then if $z:=f(\alpha_{s_{I}})$, this means that $f(x)-z$ has at least $s_{J}-s_{I}+1\geq k$ zeros (i.e. $\{\alpha_{s_{I}},\dots,\alpha_{s_{I}+(s_{J}-s_{I})}\}$ are zeros of $f(x)-z$) and it has degree at most $k-1$. Therefore $f(x)=z$ and so $f$ is constant. ∎
Proposition 2.
Let $k\geq 2$. If $q\geq 20k^{4}-90k^{3}+150k^{2}-106k+27$ then there exists an evaluation vector ${\bm{\alpha}}\in\mathbb{F}_{q}^{2k}$ such that the code $\text{RS}_{2k,k}({\bm{\alpha}})$ is optimal.
Proof.
We prove the statement by induction on $k$.
For $k=2$ from Lemma 4 we know that for $q\geq 7$ we can find an evaluation vector that has the desired properties.
Now assume that we have found a vector $\widetilde{{\bm{\alpha}}}=(\alpha_{1},\dots,\alpha_{2k-2})\in\mathbb{F}_{q}^{2k-2}$ for which the code $\text{RS}_{2k-2,k-1}(\widetilde{{\bm{\alpha}}})$ has optimal insdel distance 4, and thus by Lemma 4 has the property that ${f(\widetilde{{\bm{\alpha}}}_{I})}\neq{g(\widetilde{{\bm{\alpha}}}_{J})}$ for any $f\in\{f_{k-2}x^{k-2}+\dots+f_{1}x:f_{i}\in\mathbb{F}_{q}\textnormal{ for all $i\in\{1,\dots,k-3\}$},f_{k-2}\in\{0,1\}\}$ and $g\in\mathbb{F}_{q}[x]_{<k-1}$, with $f\neq g$ and any increasing sequences $I,J\in[2k-2]^{2k-3}$. In order to obtain a $k$-dimensional code with insdel distance $4$ in $\mathbb{F}_{q}$, again by Lemma 4 we need to make sure that we have ${f}({\bm{\alpha}}_{I})\neq{g}({\bm{\alpha}}_{J})$ for any $f\in\{f_{k-1}x^{k-1}+\dots+f_{1}x:f_{i}\in\mathbb{F}_{q}\textnormal{ for all $i\in\{1,\dots,k-2\}$},f_{k-2}\in\{0,1\}\}$ and $g\in\mathbb{F}_{q}[x]_{<k}$ with $f\neq g$ and any increasing sequences $I,J\in[2k]^{2k-1}$. We count the tuples $(\alpha_{2k-1},\alpha_{2k})\in\mathbb{F}_{q}^{2}$ that we need to exclude in order to make sure that the evaluation points $(\alpha_{1},\dots,\alpha_{2k})$ give a RS code whose largest common subsequence is of length $12k-$ 1.
Let $I=(I_{1},\dots,I_{2k-1}),J=(J_{1},\dots,J_{2k-1})\in[2k]^{2k-1}$ be increasing sequences, and denote $I^{\star}:=(I_{1},\dots,I_{2k-3})$ and $J^{\star}=(J_{1},\dots,J_{2k-3})$. We clearly have $I^{\star},J^{\star}\in[2k-2]^{2k-3}$ and there are a total of $\binom{2k-2}{2}=(2k-2)(2k-3)/2$ choices for $I^{\star}$ and $J^{\star}$ (such that they are not the same). Note that if $f({{\bm{\alpha}}}_{I})=g({{\bm{\alpha}}}_{J})$ then also $f({{\bm{\alpha}}}_{I^{\star}})=g({{\bm{\alpha}}}_{J^{\star}})$ for some $f,g\in\mathbb{F}_{q}[x]_{<k}$ with $f\neq g$. From imposing that $f({{\bm{\alpha}}}_{I^{\star}})=g({{\bm{\alpha}}}_{J^{\star}})$, we obtain the following system of $32k-$ 3 linear equations (in the variables $f_{0},\dots,f_{k-1},g_{0},\dots,g_{k-1}$):
$\begin{cases}f_{k-1}\alpha_{I_{1}}^{k-1}+\dots+f_{0}&=g_{k-1}\alpha_{J_{1}}^{k-1}+\dots+g_{0}\\ f_{k-1}\alpha_{I_{2}}^{k-1}+\dots+f_{0}&=g_{k-1}\alpha_{J_{2}}^{k-1}+\dots+g_{0}\\ &\vdots\\ f_{k-1}\alpha_{I_{2k-3}}^{k-1}+\dots+f_{0}&=g_{k-1}\alpha_{J_{2k-3}}^{k-1}+\dots+g_{0}.\end{cases}$
Now suppose that $f_{k-1}=g_{k-1}=0$. Then $f,g\in\mathbb{F}_{q}[x]_{<k-1}$, $f\neq g$, and we have $f({{\bm{\alpha}}}_{I^{\star}})=g({{\bm{\alpha}}}_{J^{\star}})$, contradicting the fact that $\widetilde{{\bm{\alpha}}}=(\alpha_{1},\dots,\alpha_{2k-3})$ was chosen such that $\text{RS}_{2k-2,k-1}(\widetilde{{\bm{\alpha}}})$ has optimal insdel distance 4. Therefore at least one of $f_{k-1}$ or $g_{k-1}$ has to be non-zero. W.l.o.g. assume $f_{k-1}\neq 0$, and we rewrite the set of equations as follows (subtracting from both sides $f_{0}$, dividing both sides by $f_{k-1}$, and rearranging):
$\begin{cases}\sum_{j=0}^{k-2}\widetilde{g}_{j}\alpha_{J_{1}}^{j}-\sum_{i=1}^{k-2}\widetilde{f}_{i}\alpha_{J_{1}}^{i}&=\alpha_{I_{1}}^{k-1}-\widetilde{g}_{k-1}\alpha_{J_{1}}^{k-1}\\ \sum_{j=0}^{k-2}\widetilde{g}_{j}\alpha_{J_{2}}^{j}-\sum_{i=1}^{k-2}\widetilde{f}_{i}\alpha_{J_{2}}^{i}&=\alpha_{I_{2}}^{k-1}-\widetilde{g}_{k-1}\alpha_{J_{2}}^{k-1}\\ &\vdots\\ \sum_{j=0}^{k-2}\widetilde{g}_{j}\alpha_{J_{2k-3}}^{j}-\sum_{i=1}^{k-2}\widetilde{f}_{i}\alpha_{J_{2k-3}}^{i}&=\alpha_{I_{2k-3}}^{k-1}-\widetilde{g}_{k-1}\alpha_{J_{2k-3}}^{k-1}\\ \end{cases}$ (5)
where $\widetilde{g}(x):=\frac{g(x)-f_{0}}{f_{k-1}}$ and $\widetilde{f}(x):=\frac{f(x)-f_{0}}{f_{k-1}}$). This system of $32k-$ 3 linear equations has $22k-$ 2 unknowns, and so we assign a fixed value in $\mathbb{F}_{q}$ to $\widetilde{g}_{k-1}$. We want to show now that there is at most one solution to this system of equations, and so $\widetilde{f}$ and $\widetilde{g}$ are uniquely determined (up to choosing $\widetilde{g}_{k-1}\in\mathbb{F}_{q}$ freely). In order to do so, it is enough to show that the kernel of the following matrix is trivial:
$M:=\begin{pmatrix}1&\alpha_{J_{1}}&\ldots&\alpha_{J_{1}}^{k-2}&\alpha_{I_{1}}&\ldots&\alpha_{I_{1}}^{k-2}\\ 1&\alpha_{J_{2}}&\ldots&\alpha_{J_{2}}^{k-2}&\alpha_{I_{2}}&\ldots&\alpha_{I_{2}}^{k-2}\\ \vdots&\vdots&\ldots&\vdots&\vdots&\ldots&\vdots\\ 1&\alpha_{J_{2k-3}}&\ldots&\alpha_{J_{2k-3}}^{k-2}&\alpha_{I_{2k-3}}&\ldots&\alpha_{I_{2k-3}}^{k-2}\\ \end{pmatrix}.$
From Lemma 2 we know that the only possible vectors in the kernel of $M$ are of the form $(0,h_{1},\dots,h_{k-2},-h_{1},\dots,-h_{k-2})$ for some $h_{1},\dots,h_{k-2}\in\mathbb{F}_{q}$. Let $(0,h_{1},\dots,h_{k-2},-h_{1},\dots,-h_{k-2})$ be in the kernel of $M$ and define $h(x)=h_{1}x+h_{2}x^{2}+\dots+h_{k-2}x^{k-2}\in\mathbb{F}[x]_{<k-1}$. We have $h({{\bm{\alpha}}}_{I})=h({{\bm{\alpha}}}_{I^{\star}})=h({{\bm{\alpha}}}_{J^{\star}})=h({{\bm{\alpha}}}_{J})$. Note also, since $I,J\in[2k]^{2k-1}$ with $d_{H}(I,J)\geq k$, we have $d_{H}(I^{\star},J^{\star})\geq k-2$. Therefore, by Claim 3 we obtain that $h$ is constant, and since $h(x)=h_{1}x+h_{2}x^{2}+\dots+h_{k-2}x^{k-2}$, this implies $h=0$.
For $I,J\in[2k]^{2k-1}$ we have $2k-2\leq I_{2k-2}<I_{2k-1}\leq 2k$ and $2k-2\leq J_{2k-2}<J_{2k-1}\leq 2k$. There are a total of $\binom{3}{2}^{2}$ options for $(I_{2k-2},I_{2k-1},J_{2k-2},J_{2k-1})$. However, we cannot have $I_{2k-2}=J_{2k-2}=2k-2$, because this would force $(I_{1},\dots,I_{2k-2})=(J_{1},\dots,J_{2k-2})$ and since $d_{H}(I,J)\geq k$ this cannot hold. Therefore there are $\binom{3}{2}^{2}-\binom{2}{1}^{2}=5$ possible options for $(I_{2k-2},I_{2k-1},J_{2k-2},J_{2k-1})$. For each of possible realization of $(I_{2k-2},I_{2k-1},J_{2k-2},J_{2k-1})$ we obtain a set of equations in $\alpha_{2k-1},\alpha_{2k}$, and so in total we have the following 5 sets of equations in $\alpha_{2k-1},\alpha_{2k}$ that would lead to problems:
$\begin{cases}&\widetilde{f}(\alpha_{2k-2})=\widetilde{g}(\alpha_{2k-1})\\ &\widetilde{f}(\alpha_{2k-1})=\widetilde{g}(\alpha_{2k})\end{cases}\begin{cases}&\widetilde{f}(\alpha_{2k-2})=\widetilde{g}(\alpha_{2k-1})\\ &\widetilde{f}(\alpha_{2k})=\widetilde{g}(\alpha_{2k})\end{cases}$
$\begin{cases}&\widetilde{f}(\alpha_{2k-1})=\widetilde{g}(\alpha_{2k-2})\\ &\widetilde{f}(\alpha_{2k})=\widetilde{g}(\alpha_{2k-1})\end{cases}\begin{cases}&\widetilde{f}(\alpha_{2k-1})=\widetilde{g}(\alpha_{2k-2})\\ &\widetilde{f}(\alpha_{2k})=\widetilde{g}(\alpha_{2k})\end{cases}$ (6)
$\begin{cases}&\widetilde{f}(\alpha_{2k-1})=\widetilde{g}(\alpha_{2k-1})\\ &\widetilde{f}(\alpha_{2k})=\widetilde{g}(\alpha_{2k}).\end{cases}$
Each of these systems has at most $(k-1)^{2}$ solutions for $(\alpha_{2k-1},\alpha_{2k})$ (because $\widetilde{f}(x),\widetilde{g}(x)\in\mathbb{F}_{q}[x]_{<k}$). Since we chose $\widetilde{g}_{k-1}\in\mathbb{F}_{q}$ freely, there are at most $5(k-1)^{2}q$ tuples $(\alpha_{2k-1},\alpha_{2k})$ that would cause problems. We also need to impose that $\alpha_{2k-1},\alpha_{2k}\notin\{\alpha_{1},\dots,\alpha_{2k-2}\}$. Let $\mathcal{A}:=\{\alpha_{1},\dots,\alpha_{2k-2}\}$. Therefore we have that whenever
$\binom{|\mathbb{F}_{q}\setminus\mathcal{A}|}{2}\geq(2k-2)(2k-3)5(k-1)^{2}q/2$
then we have enough elements in $\mathbb{F}_{q}$ to choose $(\alpha_{2k-1},\alpha_{2k})\in\mathbb{F}_{q}^{2}$ that neither satisfy the 5 set of equations, nor already show up as evaluation points in $\widetilde{{\bm{\alpha}}}$. Straightforward computations give the lower bound on $q$ as stated in the proposition. ∎
Theorem 2 is obtained by investigating the algorithm implied by the proof of Proposition 2.
Theorem 4.
There exists a deterministic algorithm that runs in $O(k^{4})$ time and outputs ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{2k})$ for which the the code $\textup{RS}_{2k,k}({\bm{\alpha}})$ can correct a single insdel error.
Proof.
We prove this theorem by describing explicitly the algorithm implied by the proof of Proposition 2. First, let $q\geq 100k^{4}$ be a prime power and define the $[4,2]_{q}$ RS code according to Remark 1. Specifically, define $\alpha_{1}=0,\alpha_{2}=1$ and $\alpha_{3},\alpha_{4}$ according to the constraints defined in Remark 1 to ensure that this $[4,2]_{q}$ RS code can correct a single insdel error. Then, run for $O(k)$ iteration the following step:
  1. (1)
    Run over all possible pairs of increasing sequences $I^{\star},J^{\star}\in[2k-2]^{2k-3}$, and all $\widetilde{g}_{k-1}\in\mathbb{F}_{q}$ and solve the linear system in (5) to get $\widetilde{f}$ and $\widetilde{g}$.
  2. (2)
    For each pair $\widetilde{f}$ and $\widetilde{g}$ from the previous step, find a set of bad pairs $(\alpha_{2k-1},\alpha_{2k})$ by solving (4.2).
  3. (3)
    Go over all possible pairs in $\mathbb{F}_{q}^{2}$ and find a good pair.
The first step takes at most $O(k^{5}q)$ time. Indeed, the number of pairs $I^{\star},J^{\star}$ is at most $\binom{2k-3}{2}$ and $\widetilde{g}_{k-1}$ takes values in $\mathbb{F}_{q}$. For each choice of $I^{\star},J^{\star}$, and $\widetilde{g}_{k-1}$, solving the linear system in (5) takes at most $O(k^{3})$ time.
Next, solving each equality in (4.2) can be done by computing the roots of a degree $k-1$ polynomial. This can be done in time $\textup{poly}(k,q)$.
In the final iteration step, we go over all $\binom{q}{2}$ possible pairs and find a good pair. Note that this step takes $O(q^{2})$ time and always succeeds according to Proposition 2. As $q=O(k^{4})$, the theorem follows. ∎
5. Discussion and Future Work
In this paper, we studied Reed-Solomon codes in the presence of insertion and deletion errors. Specifically, we investigated full-length Reed-Solomon codes and demonstrated that, when these codes have dimension 2, almost any ordering of the elements of $\mathbb{F}_{q}$ results in a code that can correct at least a single insertion or deletion error. Furthermore, using a probabilistic argument, we showed that if $q$ is large, there exists an ordering of $\mathbb{F}_{q}$ such that the $k$-dimensional Reed-Solomon code, with this ordering as the evaluation vector, can correct up to $q/(16k)$ insertion and deletion errors.
In the second part of the paper, we investigated Reed-Solomon codes with a rate of $1/2$. By the half-Singleton bound, such codes can correct at most a single insertion or deletion error. We provided an existence result for codes with a rate of $1/2$ by induction on the dimension $k$, proving that optimal codes—those capable of correcting the maximum number of insertion and deletion errors allowed by the half-Singleton bound—always exist when the underlying field size satisfies $q=O(k^{4})$. The proof of this result also led to a deterministic algorithm for constructing such codes, which runs in polynomial time.
Although we made progress toward a better understanding of Reed-Solomon codes in the context of insertion and deletion errors, several intriguing problems remain open. A natural question arising from our results is to better understand which orderings of the elements of the finite field $\mathbb{F}_{q}$ yield a full-length Reed-Solomon code that performs well against insertion and deletion errors. From experiments with magma, it appears that for $q\geq 13$, we can always construct a full-length 2-dimensional Reed-Solomon code that can correct strictly more insertion and deletion errors than classical errors (i.e., substitutions). However, it remains unclear to us which specific orderings yield such codes.
Moreover, the approach outlined in Section 4 does not appear to generalize in an obvious way to longer Reed-Solomon codes (with rates smaller than $1/2$). Current sufficient conditions on $q$ for the existence of such codes seem too restrictive, and we anticipate to get a better understanding on which field size is actually required for the existence of effective Reed-Solomon codes against insertion and deletion errors.